# Point-Spread-Function Engineering in MINFLUX: Optimality of Donut and Half-Moon Excitation Patterns

### v1

@@ Abstract
Maximally INFormative Luminescence eXcitation (MINFLUX) is an emergent localization strategy that reduces the number of photons required to achieve a given localization precision in comparison to camera-based approaches [10]. It relies on utilizing tailored excitation beam patterns sequentially to probe the location of a fluorophore. MINFLUX achieves single-digit nanometer isotropic localization precision in three dimensions fixed and living cells [11] and sub-millisecond time resolution for single-molecule tracking [12]. Using minimal number of photons, MINFLUX reduces photobleaching and enables long observation time, which makes it a suitable technique to study cellular dynamics at the molecular level [13, 14].
@@
Information theory and, in particular, the Cramér-Rao bound (CRB) are commonly used to quantify the precision of localization microscopy [15, 16, 17, 18, 19, 20] or in the more genereal context of optical imaging [21]. This quantity can then be used as a performance metric to optimize and discover new beam shapes in PSF engineering, e.g. for 3D localization in a wide field configuration [15]. To the best of our knowledge, there has been no previous work applying such a framework to MINFLUX.
@@
In this Letter, we introduce an iterative optimization procedure and a theoretical study to determine optimal beam configurations for MINFLUX. Our framework is based on a modular pipeline that takes as input the pupil function parameterized by Zernike coefficients, then models light propagation to the sample, and finally outputs the average CRB over a defined region of interest. The pupil function is optimized in an iterative manner by leveraging the automatic differentiation capability of PyTorch, a popular deep-learning library. We first confirm that the donut beam is an optimal excitation PSF under the constraint of a single excitation beam shape; we then show that two pairs of orthogonal half-moon beams lead to a further increase in the localization precision by a factor of two if more than one beam shape is possible, which has recently been implemented experimentally [22, 23].
@@
Figure 1: (a) Diagram of MINFLUX. Excitation and emission light are illustrated in green and orange, respectively. The fluorophore is represented by a yellow star. (b) Pipeline of the optimization problem. Solid and dashed lines with arrowheads indicate the forward and inverse processes, respectively. The quantities in the ellipses correspond to the inputs and outputs of the modules in the squares. (c) Intensity distribution of a donut beam on the $xy$-plane. The four dots in cyan represent the location of the center of the four donut beams in a standard MINFLUX experiment. Three of these dots lie on a circle with diameter $L=100$ nm (dashed yellow circle). The dotted red circle with radius $\tau=28$ nm represents the boundary of the FOV $\Gamma_{\tau}$ of optimization. (d) The CRB map (in nm) in the FOV $\Gamma_{\tau}$ of a standard MINFLUX experiment with beam setting as in (c). The labels on the horizontal and vertical axes describe the relative distance in nanometers to the center of the FOV for the CRB maps.
@@
The general principle of MINFLUX is as follows. The excitation light goes through a beam shaper such as a spatial light modulator (SLM) that modifies the pupil phase such that the resulting beam pattern has a zero of intensity at the focal plane where the sample is placed (Fig. 1 (a) and (c)). Such an isotropic shape has a characteristic low intensity at its center, thus termed the “zero” of the beam. Measurements in MINFLUX correspond to the fluorescence photon counts for different excitation beams, collected by an avalanche photodiode (APD). The counts are then used to infer the spatial location of the fluorescent target. In the original MINFLUX experiment [10], four donut beams are used and arranged into a specific spatial layout: one at the center of the object plane and the other three on a circle of diameter $L$ (Fig. 1 (c)) to obtain sufficient information for the inference step. The smaller $L$ is, the better the localization precision and the smaller the field of view (FOV) where the precision is high.
@@
Its derivation is detailed in the appendix. We omitted the spatial dependency of $\mu_{i}$ for conciseness. The CRB is the inverse of the FI matrix, with diagonal elements $\sigma_{\text{x}}^{2}$ and $\sigma_{\text{y}}^{2}$, which correspond to a lower bound on the variance of any unbiased estimator of the fluorophore position.
@@
The loss function $J(\mathbf{C})$ is composed of two terms. The first one is the sum of the averaged CRB values at all the locations within the FOV of optimization $\Gamma_{\tau}$, a disk of radius $\tau$. The second one is the sum of the intensities of all $K$ excitation PSFs within a given region of interest (ROI) $\Omega_{\text{h}}$. This regularization term promotes that the PSF energy remains inside that region. The nonnegative constant $\gamma$ controls the strength of the regularization term. This loss function is optimized over the matrix $\mathbf{C}$ of all $K$ Zernike-coefficient vectors $\mathbf{C}=[\mathbf{c}_{0},\ldots,\mathbf{c}_{K-1}]$.
The construction of the loss function follows a sequential structure and can be seen as a pipeline composed of modules, as illustrated in Fig. 1 (b), with the Zernike-coefficient matrix $\mathbf{C}$ as input and the loss value $J(\mathbf{C})$ as output. Leveraging the acyclic nature of the computational graph, we implement it in Pytorch to use its automatic differentiation functionality to obtain the gradient of the loss function, required for the gradient-based optimization of the parameters $\mathbf{C}$. The evaluation of (3) requires an explicit expression of the partial derivative of $\frac{\partial h_{i}}{\partial\theta_{j}}$, which is derived in the appendix. We use a custom fast Fourier transform to establish the optimal discretization parameters for the pupil function and the PSF. Based on the chirp Z-transform, it enables us to choose an arbitrarily small PSF pixel size at fixed computational cost. More details on this point and the optimization algorithm are provided in the appendix.
@@
In the first group, we adopt the original MINFLUX arrangement illustrated in Fig. 1 (c), expanded to seven beams to mitigate anisotropy in the CRB map. One beam is placed at the center of the field-of-view while the centers of the other six are regularly placed on a circle of diameter $L$ nm. All seven beams are constrained to have the same shape, a configuration that simplifies experimental implementation as it allows for a fixed phase mask in Fourier space while another optical element (e.g., a galvo mirror) translates the excitation pattern.
We initiated the optimization with a perturbed donut beam, employing a vortex phase mask $\kappa(k_{x},k_{y})$ and random initial coefficients $\mathbf{C}$. The results of this optimization, illustrated in Fig. 2 (b), exhibit a clear evolution towards a donut shape. Fig. 2 (e) displays the CRB line profiles and reveals a significant decrease in CRB values after optimization. Notably, the CRB values of our optimized beam closely match those of a reference donut beam with seven positions (both curves are superimposed). To verify the robustness of this result, we conducted additional trials under several initial conditions, including different starting points and levels of background noise. We also directly observed the evolution of the Zernike coefficients. These experiments, presented in the appendix, consistently converge to the donut shape across all scenarios and firmly establish it as a robust optimal solution for this MINFLUX configuration.
@@
The PSF shapes evolved significantly during optimization and converged to a configuration characterized by two approximately symmetric peaks separated by a valley of low intensities (Fig. 3 (c) and (d)). Further, we notice that the axes of the two pairs of beams are orthogonal to each other, oriented along the edges of the square formed by the beam centers. The associated phase masks of these beams are shown in Fig. 3 (g) and (h). This optimized configuration significantly improves the CRB values by a factor two (Fig. 3 (k) and (l)).
The optimal beams are termed “half-moon” beams as they can easily be created by a $\pi$-phase-shifted half-disk in the pupil plane (Fig. 3 (i) and (j)), which has been seen in [13]. These theoretical half-moon beams are very close to the optimized ones and yield slightly more uniform CRB maps compared to the optimized PSFs, as shown in Fig. 3 (l) and (m). A comparison of the CRB line profiles for the initial, optimized, and theoretical half-moon beams is shown in Fig. 3 (n). We tested the robustness of the half-moon shape across various optimization conditions, including scenarios where all four beams were allowed to have independent shapes. These results are provided in the appentix. Consistently, the half-moon configuration emerged as a stable and optimal solution.
To complement our optimization results, we discuss theoretical results in Theorem 1 and Proposition 2. We demonstrate how the donut and half-moon PSFs are optimal in a specific mathematical sense. A more detailed analysis, along with a rigorous proof of the theorem and numerical results, is presented in the appendix. Precision in MINFLUX is linked to how steep are the intensity variations near the center of the region of interest [10, 13], which leads to our study on how to maximize the norm of the gradient of the electric field.
Theorem 1 The half-moon beam maximizes the directional gradient norm of the electric field at the center of the field of view.
Proposition 2 The donut beam maximizes the angular average of the squared norm of the directional gradient of the electric field.
@@
Our theoretical results establish that the beam shapes obtained from our optimization framework maximize the aforementioned metrics, representing fundamental optima. However, it is important to note that these metrics do not fully capture the complexity of the MINFLUX estimation problem.
To conclude, our study leads to two key findings in the PSF engineering of MINFLUX. The donut beam is optimal when the excitation pattern is constrained to take a single shape, while the half-moon beams yield lower CRB when multiple shapes are available. These results, obtained through our PSF-engineering pipeline, are further supported by a theoretical investigation of the maximization of the gradient of the electric field. Our theoretical framework is a fast and economic approach to discover and investigate novel beam shapes for MINFLUX. Additionally, this flexible framework can be extended to optimize PSFs for 3D localization using vectorial models of the electric field propagation, potentially expanding the capabilities of MINFLUX across diverse imaging scenarios.
@@ A.1 Scalar Excitation PSF model via Zernike Polynomials
We use a scalar model to model beam propagation:
$h(x,y)=|\mathcal{F}^{-1}\left\{P(k_{\text{x}},k_{\text{y}})\right\}(x,y)|^{2}.$ (5)
(5) describes the relation between the pupil function $P$ in the Fourier space and the resulting PSF $h$ at the focal plane in the object space. The wave vector in the Fourier plane is $(k_{\text{x}},k_{\text{y}})$. The pupil function $P:\mathbb{R}^{2}\to\mathbb{C}$ is further defined as [15]
@@
The series of Zernike coefficients is represented by a $D$-dimesional vector $\mathbf{c}=(c_{0},\ldots,c_{d},\ldots,c_{D-1})$, where $c_{d}\in\mathbb{R}$, $\left\{\varphi_{d}(\cdot,\cdot)\right\}_{d=0}^{D-1}$ are the first $D$ Zernike polynomials and $\varphi_{d}:\mathbb{R}^{2}\to\mathbb{R}$. An extra phase factor $\kappa(k_{\text{x}},k_{\text{y}})$ is added to the phase. The complex term in the exponential in (6) represents a phase mask that modulates the shape of the resulting excitation PSF. When the phase mask is the ramp $\kappa(k_{\text{x}},k_{\text{y}})=\text{arctan}\left(\frac{k_{\text{x}}}{k_{\text{y}}}\right)$, the resulting PSF is a donut beam.
@@ A.2 Statistical Model of the Photon Detection in MINFLUX
The constant $\alpha$ depends on the photophysics of the emitter, the detection efficiency of the system, and the energy delivered to the sample, and $b$ is a background term.
@@ A.4 Explicit Expression of the Fisher Information matrix
The derivative of $h$ w.r.t. $x$ at location $(x-t_{x},y-t_{y})$ is
@@
Combining (24) and (23), we have an analytical expression for $\frac{\partial h}{\partial x}(x-t_{x},y-t_{y})$. Analogously, the derivative of $h$ w.r.t. $y$ at location $(x-t_{x},y-t_{y})$, $\frac{\partial h}{\partial y}(x-t_{x},y-t_{y})$, can be obtained by replacing $\mathrm{j}k_{x}$ with $\mathrm{j}k_{y}$ in the above expression. Finally, the partial derivatives of $\mu_{i}$ are easily derived using (8).
@@ A.5 Definition of the Optimization Problem
We are interested in the optimal shape of the excitation beams and look for a set of Zernike coefficients that would lead to a minimal CRB under certain experimental setups. We define such a quest as the following minimization problem of a loss function $J$
$\min_{\mathbf{C}}J(\mathbf{C})=\min_{\mathbf{C}}\left\{\sum_{(x,y)\in\Gamma}\nu(x,y)-\gamma\sum_{(x,y)\in\Omega_{\text{h}}}\left(\sum_{i=0}^{K-1}h_{i}(x,y)\right)\right\}.$ (25)
The columns in the $(D\times K)$ matrix $\mathbf{C}=[\mathbf{c}_{0},\ldots,\mathbf{c}_{K-1}]$ contain the Zernike coefficient vectors of all $K$ beams. The loss is composed of two terms. The relation of $\mathbf{c}$ and $h_{i}$ is provided in (1) and (6). The first term in $J(\mathbf{C})$ is a sum of $\nu(x,y)$, defined as the arithmetic average of the variance terms $\sigma_{\text{xx}}^{2}$ and $\sigma_{\text{yy}}^{2}$
@@ B.1 Optimization Algorithm
Our pipeline has a practical multiresolution feature. Because the input of the pipeline is a Zernike coefficient matrix that depends neither on the numerical size of the pupil nor on the PSF, we are able to run the optimization on a coarse meshgrid to save computational time and generate high-resolution images using the optimized coefficients. For optimization, the numerical sizes of the pupil and PSF are $(256\times 256)$ px and $(512\times 512)$ px, respectively. This increases computational speed while retaining sufficient sampling of the pupil function and the CRB map. After optimization, we run an extra round of the full forward pipeline that takes optimized Zernike coefficients and a larger number of sampling points. All the images in both the manuscript and the appendix are of size $(2048\times 2048)$ px unless specified otherwise.
The computational time and the evolution of the loss of each optimization experiment presented in the manuscript and the appendix are shown in Fig. 10.

### v2

@@ Abstract
Maximally INFormative Luminescence eXcitation (MINFLUX) is an emergent localization strategy that reduces the number of photons required to achieve a given localization precision in comparison to camera-based approaches [10]. It relies on utilizing tailored excitation beam patterns sequentially to probe the location of a fluorophore. MINFLUX achieves single-digit nanometer isotropic localization precision in three dimensions of fixed and living cells [11] and sub-millisecond time resolution for single-molecule tracking [12]. Using minimal number of photons, MINFLUX reduces photobleaching and enables long observation time, which makes it a suitable technique to study cellular dynamics at the molecular level [13, 14].
@@
Information theory and, in particular, the Cramér-Rao bound (CRB) are commonly used to quantify the precision of localization microscopy [15, 16, 17, 18, 19, 20] or in the more general context of optical imaging [21]. This quantity can then be used as a performance metric to optimize and discover new beam shapes in PSF engineering, e.g. for 3D localization in a wide field configuration [15]. To the best of our knowledge, there has been no previous work applying such a framework to MINFLUX.
@@
In this Letter, we introduce an iterative optimization procedure and a theoretical study to determine optimal beam configurations for MINFLUX. Our framework is based on a modular pipeline that takes as input the pupil function parameterized by Zernike coefficients, then models light propagation to the sample, and finally outputs the average CRB over a defined region of interest. The pupil function is optimized by an iterative gradient-descent-type algorithm, leveraging the automatic differentiation capability of PyTorch, a popular deep-learning library. We first confirm that the donut beam is an optimal excitation PSF under the constraint of a single excitation beam shape; we then show that two pairs of orthogonal half-moon beams lead to a further increase in the localization precision by a factor of two if more than one beam shape is possible, which has recently been implemented experimentally [22, 23].
@@
Figure 1: (a) Diagram of MINFLUX. Excitation and emission light are illustrated in green and orange, respectively. The fluorophore is represented by a yellow star. (b) Intensity distribution of a donut beam on the $xy$-plane. The four dots in cyan represent the location of the center of the four donut beams in a standard MINFLUX experiment. Three of these dots lie on a circle with diameter $L=100$ nm (dashed yellow circle). The dotted red circle with radius $\tau=28$ nm represents the boundary of the FOV $\Gamma_{\tau}$ of optimization. (c) Pipeline of the optimization problem. Solid and dashed lines with arrowheads indicate the forward and inverse processes, respectively. The quantities in the ellipses correspond to the inputs and outputs of the modules in the squares. (d) The CRB map (in nm) in the FOV $\Gamma_{\tau}$ of a standard MINFLUX experiment with beam setting as in (b) for 100 photons. The labels on the horizontal and vertical axes describe the relative distance in nanometers to the center of the FOV for the CRB maps.
@@
The general principle of MINFLUX is as follows. The excitation light goes through a beam shaper such as a spatial light modulator (SLM) that modifies the pupil phase such that the resulting beam pattern has a zero of intensity at the focal plane where the sample is placed (Fig. 1 (a) and (b)). Such an isotropic shape has a characteristic low intensity at its center, thus termed the “zero” of the beam. Measurements in MINFLUX correspond to the fluorescence photon counts for different excitation beams, collected by an avalanche photodiode (APD). The counts are then used to infer the spatial location of the fluorescent target. In the original MINFLUX experiment [10], four donut beams are used and arranged into a specific spatial layout: one at the center of the object plane and the other three on a circle of diameter $L$ (Fig. 1 (b)) to obtain sufficient information for the inference step. The smaller $L$ is, the better the localization precision and the smaller the field of view (FOV) where the precision is high.
@@
Its derivation is detailed in Section 1.C of the Supplement. We omitted the spatial dependency of $\mu_{i}$ for conciseness. The CRB is the inverse of the FI matrix, with diagonal elements $\sigma_{\text{x}}^{2}$ and $\sigma_{\text{y}}^{2}$, which correspond to a lower bound on the variance of any unbiased estimator of the fluorophore position.
@@
The loss function $J(\mathbf{C})$ is composed of two terms. The first one is the sum of the averaged CRB values at all the locations within the FOV of optimization $\Gamma_{\tau}$, a disk of radius $\tau$. The second one is the sum of the intensities of all $K$ excitation PSFs within a given region of interest (ROI) $\Omega_{\text{h}}$. This regularization term encourages that the PSF energy remains inside that region. The nonnegative constant $\gamma$ controls the strength of the regularization term. This loss function is optimized over the matrix $\mathbf{C}$ of all $K$ Zernike-coefficient vectors $\mathbf{C}=[\mathbf{c}_{0},\ldots,\mathbf{c}_{K-1}]$.
The construction of the loss function follows a sequential structure and can be seen as a pipeline composed of modules, as illustrated in Fig. 1 (c), with the Zernike-coefficient matrix $\mathbf{C}$ as input and the loss value $J(\mathbf{C})$ as output. Leveraging the acyclic nature of the computational graph, we implement it in Pytorch to use its automatic differentiation functionality to obtain the gradient of the loss function, required for the gradient-based optimization of the parameters $\mathbf{C}$. The evaluation of (3) requires an explicit expression of the partial derivative of $\frac{\partial h_{i}}{\partial\theta_{j}}$, which is derived in Section 1.E of the Supplement. We use a custom fast Fourier transform to establish the optimal discretization parameters for the pupil function and the PSF. Based on the chirp Z-transform, it enables us to choose an arbitrarily small PSF pixel size at fixed computational cost. More details on this point and the optimization algorithm are provided in Section 2 of the Supplement.
@@
In the first group, we adopt the original MINFLUX arrangement illustrated in Fig. 1 (b), expanded to seven beams to mitigate anisotropy in the CRB map. One beam is placed at the center of the field-of-view while the centers of the other six are regularly placed on a circle of diameter $L$ nm. All seven beams are constrained to have the same shape, a configuration that simplifies experimental implementation as it allows for a fixed phase mask in Fourier space while another optical element (e.g., a galvo mirror) translates the excitation pattern.
We initiated the optimization with a perturbed donut beam, employing a vortex phase mask $\kappa(k_{x},k_{y})$ and random initial coefficients $\mathbf{C}$. The results of this optimization, illustrated in Fig. 2 (b), exhibit a clear evolution towards a donut shape. Fig. 2 (e) displays the CRB line profiles and reveals a significant decrease in CRB values after optimization. Notably, the CRB values of our optimized beam closely match those of a reference donut beam with seven positions (both curves are superimposed). To verify the robustness of this result, we conducted additional trials under several initial conditions, including different starting points and levels of background noise. We also directly observed the evolution of the Zernike coefficients. These experiments, presented in the Supplement, consistently converge to the donut shape across all scenarios and firmly establish it as a robust optimal solution for this MINFLUX configuration.
@@
The PSF shapes evolved significantly during optimization and converged to a configuration characterized by two approximately symmetric peaks separated by a valley of low intensities (Fig. 3 (c) and (d)). Further, we notice that the axes of the two pairs of beams are orthogonal to each other, oriented along the edges of the square formed by the beam centers. The associated phase masks of these beams are shown in Fig. 3 (g) and (h). This optimized configuration significantly improves the CRB values by a factor of two (Fig. 3 (k) and (l)).
The optimal beams resemble the existing “half-moon” beams which have been seen in [13]. In practice, these shapes can be generated by a $\pi$-phase-shifted half-disk in the pupil plane (Fig. 3 (i) and (j)). These theoretical half-moon beams yield slightly more uniform CRB maps compared to the optimized PSFs, as shown in Fig. 3 (l) and (m). A comparison of the CRB line profiles for the initial, optimized, and theoretical half-moon beams is shown in Fig. 3 (n). We tested the robustness of the half-moon shape across various optimization conditions, including scenarios where all four beams were allowed to have independent shapes. These results are provided in the Supplement. Consistently, the half-moon configuration emerged as a stable and optimal solution.
To complement our optimization results, we discuss theoretical results in Theorem 1 and Proposition 2. We demonstrate how the donut and half-moon PSFs are optimal in a specific mathematical sense. A more detailed analysis, along with a rigorous proof of the theorem and numerical results, is presented in Section 3 in the Supplement. Precision in MINFLUX is linked to the steepness of the intensity variations near the center of the region of interest [10, 13], which leads to our study on how to maximize the norm of the gradient of the electric field.
Theorem 1 The half-moon beam maximizes the directional derivative norm of the electric field at the center of the field of view.
Proposition 2 The donut beam maximizes the angular average of the squared norm of the directional derivative of the electric field.
@@
Our theoretical results establish that the beam shapes obtained from our optimization framework maximize the aforementioned metrics, representing fundamental optima. However, they do not account for shifts and the estimation problem, which necessitates our Fisher-information framework.
To conclude, our study leads to two key findings in the PSF engineering of MINFLUX. The donut beam is optimal when the excitation pattern is constrained to take a single shape, while the half-moon beams yield lower CRB when multiple shapes are available. These results, obtained through our PSF-engineering pipeline, are further supported by a mathematical investigation of the maximization of the gradient of the electric field. Our theoretical framework is a fast and economic approach to discover and investigate novel beam shapes for MINFLUX. Additionally, this flexible framework can be extended to optimize PSFs for 3D localization using vectorial models of the electric field propagation, potentially expanding the capabilities of MINFLUX across diverse imaging scenarios.
@@ A.1 Scalar Excitation PSF model via Zernike Polynomials
We use a scalar model (1)for the beam propagation. It describes the relation between the pupil function $P$ in the Fourier space and the resulting PSF $h$ at the focal plane in the object space. The wave vector in the Fourier plane is $(k_{\text{x}},k_{\text{y}})$. The pupil function $P:\mathbb{R}^{2}\to\mathbb{C}$ is further defined as [15]
@@
The series of Zernike coefficients is represented by a $D$-dimensional vector $\mathbf{c}=(c_{0},\ldots,c_{d},\ldots,c_{D-1})$, where $c_{d}\in\mathbb{R}$, $\left\{\varphi_{d}(\cdot,\cdot)\right\}_{d=0}^{D-1}$ are the first $D$ Zernike polynomials and $\varphi_{d}:\mathbb{R}^{2}\to\mathbb{R}$. An extra phase factor $\kappa(k_{\text{x}},k_{\text{y}})$ is added to the phase. The complex term in the exponential in (5) represents a phase mask that modulates the shape of the resulting excitation PSF. When the phase mask is the ramp $\kappa(k_{\text{x}},k_{\text{y}})=\text{arctan}\left(\frac{k_{\text{x}}}{k_{\text{y}}}\right)$, the resulting PSF is a donut beam.
@@ A.2 Statistical Model of the Photon Detection in MINFLUX
The constant $\alpha$ depends on the photo-physics of the emitter, the detection efficiency of the system, and the energy delivered to the sample, and $b$ is a background term.
@@ A.4 Explicit Expression of the Fisher Information Matrix
The derivative of $h$ with respect to $x$ at location $(x-t_{x},y-t_{y})$ is
@@
Combining (23) and (22), we have an analytical expression for $\frac{\partial h}{\partial x}(x-t_{x},y-t_{y})$. Analogously, the derivative of $h$ with respect to $y$ at location $(x-t_{x},y-t_{y})$, $\frac{\partial h}{\partial y}(x-t_{x},y-t_{y})$, can be obtained by replacing $\mathrm{j}k_{x}$ with $\mathrm{j}k_{y}$ in the above expression. Finally, the partial derivatives of $\mu_{i}$ are easily derived using (7).
@@ A.5 Definition of the Optimization Problem
We are interested in the optimal shape of the excitation beams and look for a set of Zernike coefficients that would lead to a minimal CRB under certain experimental setups. We define such a quest as the following minimization problem of a loss function $J$ defined in (4). The columns in the $(D\times K)$ matrix $\mathbf{C}=[\mathbf{c}_{0},\ldots,\mathbf{c}_{K-1}]$ contain the Zernike coefficient vectors of all $K$ beams. The loss is composed of two terms. The relation of $\mathbf{c}$ and $h_{i}$ is provided in (1) and (5). The first term in $J(\mathbf{C})$ is a sum of $\nu(x,y)$, defined as the arithmetic average of the variance terms $\sigma_{\text{xx}}^{2}$ and $\sigma_{\text{yy}}^{2}$
@@ B.1 Optimization Algorithm
Our pipeline has a practical multi-resolution feature. Because the input of the pipeline is a Zernike coefficient matrix that depends neither on the numerical size of the pupil nor on the PSF, we are able to run the optimization on a coarse meshgrid to save computational time and generate high-resolution images using the optimized coefficients. For optimization, the numerical sizes of the pupil and PSF are $(256\times 256)$ px and $(512\times 512)$ px, respectively. This increases computational speed while retaining sufficient sampling of the pupil function and the CRB map. After optimization, we run an extra round of the full forward pipeline that takes optimized Zernike coefficients and a larger number of sampling points. All the images in both the manuscript and the Supplement are of size $(2048\times 2048)$ px unless specified otherwise.
The computational time and the evolution of the loss of each optimization experiment presented in the manuscript and the supplement are shown in Fig. 10.